# Leveraging Generative AI for Smart City Digital Twins: A Survey on the Autonomous Generation of Data, Scenarios, 3D City Models, and Urban Designs

### v1

@@ 2 Generative AI Models
Generative AI models are a class of artificial intelligence techniques designed to generate new data samples similar to a given set of input data (Bandi et al., 2023). These models can produce a wide range of outputs, including images, text, sound, and video, and are particularly notable for their ability to create realistic, novel, and often indistinguishable data from actual human-generated content (Eigenschink et al., 2021), allowing them to support data-driven research application in various urban science and smart city sectors. In addition, Generative AI techniques and services enabled through Large Language Models (LLMs), such as ChatGPT, have demonstrated their unique capabilities in interpreting human natural language and automating code generation and information, therefore, have been applied to develop autonomous systems in various research fields, such as geoinformatics (Li and Ning, 2023), chemical research (Boiko et al., 2023), and transportation (Cui et al., 2024). We discovered the previously published research articles through the Scopus database. We summarized popular generative AI models used in major urban digital twin application areas (as identified in the motivation section). Based on our search combined with summaries proposed by other review articles (Lin et al., 2023; Bandi et al., 2023), this survey paper primarily targets four generative AI models, including the Generative Adversarial Networks, Variational Autoencoders, ransformer-Based Models, and Generative Diffusion Models. These models and their urban applications are briefly and individually discussed in the following subsections.
@@ 2.1 Generative Adversarial Networks
Generative Adversarial Networks (GANs) represent a pivotal innovation in the field of artificial intelligence, particularly within the domain of deep learning (Pan et al., 2019b; Goodfellow et al., 2014). These models introduce a novel framework for generative modeling, leveraging the power of two neural networks—a generator and a discriminator—in a competitive setting. The generator aims to produce data indistinguishable from real data, while the discriminator endeavors to distinguish between real and generated data (Goodfellow et al., 2020; Creswell et al., 2018). This adversarial process results in the generator creating increasingly realistic outputs. Over time, GANs have evolved into various forms, such as Conditional GANs, which generate data based on specific conditions (Mirza and Osindero, 2014), and CycleGANs, used for style transfer tasks (Zhu et al., 2017). Their vast and transformative applications range from image and video generation to drug discovery and advanced image editing. GANs have revolutionized how machines understand and recreate complex data patterns, making them a cornerstone technology in modern AI research and development. Their continued evolution promises further breakthroughs in AI’s capability to generate realistic, nuanced, and context-aware content.
@@
    Data Augmentation: the process is used to increase the size and diversity of an existing dataset. GANs can create additional data with improved spatiotemporal resolution for training predictive models, enhancing the robustness of these models without the need to collect more real-world data. In recent years, GANs have been applied to augment urban data to improve the prediction accuracy of many urban variables in areas that include transportation management (Dabboussi, 2023; Lin et al., 2023), building energy optimization (Fan et al., 2022), and infrastructure management (Gwon et al., 2023).
@@
    Synthetic Data for Scenario Generation: the primary goal of synthetic data generation is to create entirely new data points that are not present in the original dataset. This is especially useful when the amount of real data available is limited (e.g., hypothetical scenarios or corner cases) or when using real data is restricted due to privacy concerns (e.g., personal or sensitive information). GANs are employed to generate synthetic data to define many hypothetical scenarios (Sankar, 2020) with a wide range of urban management interests, such as promoting transportation planning (Wu et al., 2020; Demetriou et al., 2020; Rempe et al., 2022), urban mobility management (Wang et al., 2020), power system operations (Yuan et al., 2021; Huang et al., 2023; Rizzato et al., 2022), building energy modeling (Wang and Hong, 2020; Zhang and Zhang, 2019), and water management (Koochali et al., ).
@@ 2.1.2 CycleGANs
CycleGAN, short for Cycle-Consistent Generative Adversarial Networks, represents a significant advancement in the field of machine learning and computer vision. This innovative framework facilitates image-to-image translation tasks without requiring paired training data (Zhu et al., 2017). This is a notable departure from traditional methods that rely heavily on paired datasets, which are often difficult and expensive to obtain. CycleGAN leverages a novel approach that uses two sets of Generative Adversarial Networks (GANs) and a cycle-consistency loss function CycleGAN. This ensures that the inputs closely resemble the original inputs when translated to another domain and then back. This breakthrough has wide-ranging applications, from photo-realistic image translation to style transfer and even in tasks such as object transfiguration and season transfer in images. Its ability to learn mappings between different domains without paired examples makes CycleGAN a powerful tool in unsupervised learning, offering new possibilities in various smart city fields such as autonomous driving (Rajagopal et al., 2023), augmented reality and virtual reality (Kim et al., 2021), and context-specific data augmentation (Zhou et al., 2023a; Li et al., 2021).
@@ 2.2 Variational Autoencoders
    Missing Data Imputation: VAEs can facilitate the process of replacing missing data in a dataset with substituted values. The goal is to address gaps and poor quality in data to allow for effective analysis and decision-making. Many previous urban science applications employed the VAE to support data imputation and enrichment in the interest of improving the prediction and forecasting in traffic (Chen et al., 2021; Boquet et al., 2020; Pereira et al., 2020), energy (Shen et al., 2021; Liu and Zhang, 2021), and urban water resources management (Zhang et al., 2023; Wang et al., 2023b; Xie et al., 2019).
@@
2.3 Transformer-Based Models
Transformer-based models, a groundbreaking innovation in natural language processing (NLP), have revolutionized how machines understand and generate human language (Gillioz et al., 2020). Characterized by their unique architecture, which relies on self-attention mechanisms, transformers facilitate the processing of data sequences, such as text, in a manner that captures contextual relationships more effectively than prior methods (Vaswani et al., 2017). This architecture has been instrumental in the development of advanced language models like ChatGPT, which demonstrates exceptional capabilities in generating human-like text, answering questions, and even assisting in coding and analytics tasks (Feng et al., 2023b; Roumeliotis and Tselikas, 2023).
In recent years, the scientific research community has shown a growing interest in transformer-based models, especially the chatGPT, primarily due to their versatility and efficiency in understanding and contextualizing text for data and knowledge discovery, as well as their intelligent capability for generating codes in multiple programming languages (Sohail et al., 2023). In the urban science sector, GPT and its Application Programming Interface (API) have been used to support a variety of innovative applications in the following aspects:
@@ 2.3 Transformer-Based Models
    Autonomous Expert System: leveraging the transformer-based models’ capability to automate domain-specific data discovery (Li and Zhang, 2023) and code generation (Li and Ning, 2023; Feng et al., 2023b), and facilitate decision-support (Singh, 2023; Tan et al., 2023) for complex environmental and urban problems.
@@ 2.4 Generative Diffusion Models
Diffusion Models, a class of generative models that have gained prominence in recent years, represent a novel approach in the field of machine learning for synthesizing high-quality data (Ho et al., 2020; Sohl-Dickstein et al., 2015). These models operate on the principle of gradually transforming a distribution of random noise into a distribution of structured data by mimicking the process of diffusion in physics and have broken the long-time dominance of GANs (Yang et al., 2023). The roots of Diffusion Models can be traced back to earlier works in denoising and score-based generative models. Still, the research in the late 2010s truly harnessed their potential in generating complex data such as images and audio. In scientific research, diffusion Models hold significant promise due to their ability to generate realistic and varied data samples, improve likelihood estimation, and handle data with special structures (Yang et al., 2023; Cao et al., 2022). In the urban science and management sector, the recently emerging generative diffusion models are increasingly applied in smart applications to facilitate the simulation and modeling of urban dynamics and processes (Zhou et al., 2023b; Zhong et al., 2023), as well as generative urban design and optimization (Du et al., 2023; Su et al., 2023; He et al., 2023).
@@
3 Generative AI in Smart City Applications
@@
This section will provide a structured review of how the recent advancements in generative AI are revolutionizing key areas of urban science research and smart city applications, including traffic and transportation management, mobility research, building operations, energy management, and infrastructure planning. We primarily focus on the applications of the Generative AI models, including GANs, VAEs, transformer-based models, and generative diffusion models. The following subsections will delve into specific application areas of Generative AI models in each of the mentioned smart city areas (as depicted in Figure 2) to highlight their transformative ability to address challenges in urban digital twin applications.
@@
Figure 2: Generative AI application areas in major smart city sectors.
3.1 Data Augmentation
Data augmentation refers to techniques used to increase the amount of data by adding slightly modified copies of already existing data or newly created synthetic data from existing data (Lemley et al., 2017). This is particularly common in machine learning and deep learning, where more training data leads to better model performance. Data augmentation is primarily about creating additional data points to improve the robustness and accuracy of models in tasks like image recognition, classification, and predictions (Moreno-Barea et al., 2020). Generative Adversarial Networks (GANs) and Variational Autoencoders (VAEs) are prevalent in smart city applications for data augmentation, given their proficiency in generating high-quality synthetic data. However, generative diffusion models, known for their strengths in denoising and data augmentation, are primarily utilized in medical research rather than urban science. This suggests a potential untapped area for these models in urban applications, where their unique capabilities could offer novel solutions to complex urban challenges.
3.1.1 Mobility and Transportation Management
Generative AI plays a crucial role in transforming transportation and urban mobility management by leveraging data augmentation techniques. It enhances the accuracy and reliability of predictive models, enabling more efficient traffic flow, route optimization, and smart city planning by generating realistic, varied datasets that mimic complex urban environments.

### v2

@@ 2 Generative AI Models
Generative AI models are a class of artificial intelligence techniques designed to generate new data samples similar to a given set of input data (Bandi et al., 2023). These models can produce a wide range of outputs, including images, text, sound, and video, and are particularly notable for their ability to create realistic, novel, and often indistinguishable data from actual human-generated content (Eigenschink et al., 2021), allowing them to support data-driven research application in various urban science and smart city sectors. In addition, Generative AI techniques and services enabled through Large Language Models (LLMs), such as ChatGPT, have demonstrated their unique capabilities in interpreting human natural language and automating code generation and information, therefore, have been applied to develop autonomous systems in various research fields, such as geoinformatics (Li and Ning, 2023), chemical research (Boiko et al., 2023), and transportation (Cui et al., 2024a).
Based on our search, combined with summaries proposed by other review articles (Lin et al., 2023b; Bandi et al., 2023), and the scope of GenAI models defined by Bengesi et al. (2024), this paper primarily targets four GenAI models: Generative Adversarial Networks, Variational Autoencoders, Generative Pre-trained Transformers, and Generative Diffusion Models. These models and their urban applications are briefly and individually discussed in the following subsections.
@@ 2.1 Generative Adversarial Networks
Generative Adversarial Networks (GANs) represent a pivotal innovation in the field of artificial intelligence, particularly within the domain of deep learning (Pan et al., 2019b; Goodfellow et al., 2014). These models introduce a novel framework for generative modeling, leveraging the power of two neural networks—a generator and a discriminator—in a competitive setting. The generator aims to produce data indistinguishable from real data, while the discriminator endeavors to distinguish between real and generated data (Goodfellow et al., 2020; Creswell et al., 2018). This adversarial process results in the generator creating increasingly realistic outputs. Over time, GANs have evolved into various forms, such as Conditional GANs, which generate data based on specific conditions (Mirza and Osindero, 2014), and CycleGANs, used for style transfer tasks (Zhu et al., 2017). Their vast and transformative applications range from image and video generation to drug discovery and advanced image editing. GANs have revolutionized how machines understand and recreate complex data patterns, making them a cornerstone technology in modern AI research and development. Their continued evolution promises further breakthroughs in AI’s capability to generate realistic content based on specific contexts.
@@
    Data Augmentation: the process is used to increase the size and diversity of an existing dataset. GANs can create additional data with improved spatiotemporal resolution for training predictive models, enhancing the robustness of these models without the need to collect more real-world data. In recent years, GANs have been applied to augment urban data to improve the prediction accuracy of many urban variables in areas that include transportation management (Dabboussi, 2023; Lin et al., 2023b), building energy optimization (Fan et al., 2022b), and infrastructure management (Gwon et al., 2023).
@@
    Synthetic Data for Scenario Generation: the primary goal of synthetic data generation is to create entirely new data points that are not present in the original dataset. This is especially useful when the amount of real data available is limited (e.g., hypothetical scenarios or corner cases) or when using real data is restricted due to privacy concerns (e.g., personal or sensitive information). GANs are employed to generate synthetic data to define many hypothetical scenarios (Sankar, 2020) with a wide range of urban management interests, such as promoting transportation planning (Wu et al., 2020; Demetriou et al., 2020; Rempe et al., 2022), urban mobility management (Wang et al., 2020a), power system operations (Yuan et al., 2021; Huang et al., 2023b; Rizzato et al., 2022), building energy modeling (Wang and Hong, 2020; Zhang and Zhang, 2019), and water management (Koochali et al., 2023).
@@ 2.1.2 CycleGANs
CycleGAN, short for Cycle-Consistent Generative Adversarial Networks, represents a significant advancement in the field of machine learning and computer vision. This innovative framework facilitates image-to-image translation tasks without requiring paired training data (Zhu et al., 2017). This is a notable departure from traditional methods that rely heavily on paired datasets, which are often difficult and expensive to obtain. CycleGAN leverages a novel approach that uses two sets of Generative Adversarial Networks (GANs) and a cycle-consistency loss function CycleGAN. This ensures that the inputs closely resemble the original inputs when translated to another domain and then back. This breakthrough has wide-ranging applications, from photo-realistic image translation to style transfer and even in tasks such as object transfiguration and season transfer in images. Its ability to learn mappings between different domains without paired examples makes CycleGAN a powerful tool in unsupervised learning, offering new possibilities in various smart city fields such as autonomous driving (Rajagopal et al., 2023), augmented reality and virtual reality (Kim et al., 2021), and context-specific data augmentation (Zhou et al., 2023a; Li et al., 2021b).
2.1.3 Comparison between GANs
Building upon previous literature that has provided comprehensive reviews of various GAN variants and their applications in different urban science sectors, including transportation (Lin et al., 2023b) and the built environment (Wu et al., 2022a), this subsection compares different GANs through Table 2. The table summarizes the differences between various types of GANs, highlighting their strengths and weaknesses, and describing their popular applications.
Table 2: Comparison between GAN and its variants.
Model Rationale Strengths Weaknesses Popular Applications
GANs Two neural networks (Generator & Discriminator) compete against each other. Produce high-quality, realistic samples (Shmelkov et al., 2018). Mode collapse and training instability (Thanh-Tung and Tran, 2020). Image, vector, and raster data generation, style transfer, data augmentation (Wu et al., 2022a).
cGANs GANs with additional information (labels) provided to both generator and discriminator. Generate class-specific samples, improved control over output (Rangwani et al., 2022). Requires labeled data, potential for overfitting (Gauthier, 2014). Synthetic and conditional data generation, simulation scenario generation (Durgadevi et al., 2021).
CycleGANs Use two GANs to learn mappings between two domains without paired examples. Generate high-quality sample and capture dependencies between features in the data and images (Ma, 2024). Training complexity and potential for artifacts (Ma, 2024). Aerial-to-map translations, style transfer and unsupervised image-to-image translation (Ma, 2024; Fontanini et al., 2022).
@@ 2.2 Variational Autoencoders
    Missing Data Imputation: VAEs can facilitate the process of replacing missing data in a dataset with substituted values. The goal is to address gaps and poor quality in data to allow for effective analysis and decision-making. Many previous urban science applications employed the VAE to support data imputation and enrichment in the interest of improving the prediction and forecasting in traffic (Chen et al., 2021; Boquet et al., 2020; Pereira et al., 2020), energy (Shen et al., 2021; Liu and Zhang, 2021), and urban water resources management (Zhang et al., 2023; Wang et al., 2023c; Xie et al., 2019).
@@
2.3 Generative Pre-trained Transformers
Generative pre-trained transformers (GPT) are a type of Large Language Models (LLMs), a groundbreaking innovation in natural language processing (NLP), and have revolutionized how machines understand and generate human language (Gillioz et al., 2020). Characterized by their unique architecture, which relies on self-attention mechanisms, transformers facilitate the processing of data sequences, such as text, in a manner that captures contextual relationships more effectively than prior methods (Vaswani et al., 2017). This architecture has been instrumental in the development of advanced language models like ChatGPT, which demonstrates exceptional capabilities in generating human-like text, answering questions, and even assisting in coding and analytics tasks (Feng et al., 2023b; Roumeliotis and Tselikas, 2023).
In recent years, the scientific research community has shown a growing interest in transformer-based models, especially the OpenAI’s ChatGPT and Google’s BERT, primarily due to their versatility and efficiency in understanding and contextualizing text for data and knowledge discovery, as well as their intelligent capability for generating codes in multiple programming languages (Sohail et al., 2023). In the urban science sector, GPT and its Application Programming Interface (API) have been used to support a variety of innovative applications in the following aspects:
@@ 2.3 Generative Pre-trained Transformers
    Autonomous Expert System: leveraging the GPTs’ capability to automate domain-specific data discovery (Li and Zhang, 2023) and code generation (Li and Ning, 2023; Feng et al., 2023b), ontology and knowledge graph creation (Tupayachi et al., 2024), and facilitate decision-support (Singh, 2023; Tan et al., 2023) for complex environmental and urban problems.
Early versions of Large Language Models (LLMs) and their associated applications, such as Contextual Predictive Text systems (CPTs), faced significant limitations in their knowledge base due to the static nature of the data on which they were trained (Bubeck et al., 2023). Concerns are raised regarding the factuality of AI-Generated response to domain-specific questions that require the most current and dynamic information, and specilized knowledge (Cao et al., 2021; Bender et al., 2021). To address these shortcomings, there is a need to improve the quality and accuracy of the AI-generated content with vast factual knowledge retrieved from online knowledge base (Borgeaud et al., 2022). Recent advancements in AI and computing technologies have introduced various tuning techniques for LLMs, together with the Retrieval-Augmented Generation (RAG) technique, to enhance the GPT’ ability to access and utilize up-to-date data and specialized domain knowledge base (Lewis et al., 2020), thereby improving their performance in completing tasks related to specialized domains and ensuring more updated and relevant responses with high factual accuracy (Wang et al., 2023a).
@@ 2.4 Generative Diffusion Models
Diffusion Models, a class of generative models that have gained prominence in recent years, represent a novel approach in the field of machine learning for synthesizing high-quality data (Ho et al., 2020; Sohl-Dickstein et al., 2015). These models operate on the principle of gradually transforming a distribution of random noise into a distribution of structured data by mimicking the process of diffusion in physics and have broken the long-time dominance of GANs (Yang et al., 2023). The roots of Diffusion Models can be traced back to earlier works in Denoising Diffusion Probabilistic Models (DDPMs) Ho et al. (2020) and score-based generative models (Song et al., 2020). Still, the research in the late 2010s truly harnessed their potential in generating complex data such as images and audio. In scientific research, diffusion Models hold significant promise due to their ability to generate realistic and varied data samples, improve likelihood estimation, and handle data with special structures (Yang et al., 2023; Cao et al., 2022). In the urban science and management sector, the recently emerging generative diffusion models are increasingly applied in smart applications to facilitate the simulation and modeling of urban dynamics and processes (Zhou et al., 2023b; Zhong et al., 2023), as well as generative urban design and optimization (Du et al., 2023; Su et al., 2023; He et al., 2023).
@@
2.5 Generative Model Comparison
Based on numerous literature reviews and surveys on popular GenAI models, we present a comparative summary of different generative AI models in Table 3, highlighting their strengths, weaknesses, and applications. It is worth mentioning that many generic GenAI applications, such as the generation of images, vectors, and 3D models, can be tailored to generate specialized GIS data for smart city applications. This is because raster data, shapefiles, and 3D terrains are essentially spatially referenced images, vectors, and 3D models respectively. Theoretically, the currently popular applications of GenAI models for artistic and recreational purposes can be tailored to support urban science research and smart city applications. This can be achieved by training these models using urban datasets and simulation outputs.
Table 3: Comparative summary of the strength, weakness, and applications of popular GenAI models.
Model Rationale Strengths Weaknesses Popular Application
GANs Two neural networks (Generator and Discriminator) compete against each other Produce high-quality and realistic samples (Shmelkov et al., 2018). Mode collapse and training instability (Thanh-Tung and Tran, 2020). Image, vector, and raster data generation, style transfer, data augmentation (Wu et al., 2022a).
VAEs Latent space modeling via encoder-decoder architecture with variational inference. Robust training through explicit modeling of data distribution, the usage of stochastic gradient descent, and regularization through Kullback-Leibler (KL) divergence (Cemgil et al., 2020; Kingma et al., 2017). Blurriness in generated images, less sharp results (Zheng et al., 2020). Data compression, image generation, anomaly detection, and dimension reduction (Mak et al., 2023).
GPT Attention mechanisms to capture dependencies across input sequences. Natural language understanding and generation, contextual understanding, and few-shot and zero-shot learning with minimum training data (Ray, 2023). Limited sequential modeling, high computational cost, and potential bias (Kalyan, 2023; Brown et al., 2020). Natural language processing, image generation, human language understanding, code generation, translation tasks (Kalyan, 2023)
Diffusion Models Gradually unfold the data distribution through iterative refinement. Generate high-quality samples that capture complex dependencies and may outperform GANs (Dhariwal and Nichol, 2021). Computationally intensive and sensitive to initialization (Song et al., 2020). Generation of image, audio, 3D model, and graph-structured data generation, as well as image inpainting and anomaly detection (Song et al., 2020).
3 GenAI in Smart City Applications
@@
This section will provide a structured review of how recent advancements in GenAI are revolutionizing key areas of urban science research and smart city applications. These areas include traffic and transportation management, mobility research, building operations, energy management, and infrastructure planning. We will primarily focus on the applications of various GenAI models, such as GANs, VAEs, GPTs, and generative diffusion models. Our aim is to address research question Q3 by exploring how GenAI technologies are applied in the smart city sector through comprehensive reviews of previous studies.
The following subsections will delve into specific GenAI applications in each of the aforementioned urban application areas (as depicted in Figure 3 and defined in 1). These subsections aim to highlight GenAI’s ability to address challenges in urban digital twin applications discussed in Section 1.2. Within each subsection, we will select a handful of applications to discuss in detail, focusing on their methodology and implementation strategy. This selection is based on the study’s potential for integration into urban digital twins, with a preference for data-driven applications that align with the digital twin’s primary functions of enabling situational awareness, predictive analysis, and informed decision support.
@@
Figure 3: GenAI application areas in major smart city sectors.
3.1 Mobility and Transportation Management
From the 259 articles selected for their relevance to urban transportation, we performed an additional screening of their titles, keywords, and abstracts. This allowed us to identify the primary GenAI use cases that can be integrated into urban digital twins within the context of intelligent transportation systems and smart mobility. The various use cases are detailed in the following subsections.
3.1.1 Transportation Data Augmentation